# A Diagnostic to Find and Help Combat Positivity Issues – with a Focus on Continuous Treatments

### v1

@@ 4 Simulation Study
As previously stated, publications in the field of causal research are rather focused on static interventions. Other intervention schemes and MTPs are much less represented in research. Therefore, MTPs are not well established and not generally well understood by practitioners. However, the use of MTPs is often recommended as a response to positivity issues, as they “have the advantage that they can be designed to satisfy the positivity assumption required for causal inference” [8, p. 846]. In the following simulation study, we provide insights into the behaviour of some of the most popular MTPs in different sparsity situations. We show that it is not easy to know when positivity issues influence an estimand and that our diagnostic bridges that gap, providing the information needed to decide which estimands are feasible and which are likely to be heavily biased due to sparsity issues. The code for this simulation study is made available in full on Github.
@@ 4.1 Methods
The simulation is kept as simple as possible on purpose in order to illustrate the issues MTPs face in light of positivity violations and how the diagnostic can help understanding the problem. It is meant as a proof of concept. An analysis of a more complex scenario is presented in the form of an applied study in Chapter 5. The data are simulated using the simple structural causal model $O=(L,A,Y)\sim P_{0}$, where $L$ is a binary confounder, $A$ the continuous treatment variable, and $Y$ the continuous outcome:
@@
The function $g_{A}(.)$ and variable $\sigma_{a}^{2}$ control the distribution of $A$ and can be used to induce problems regarding the positivity assumption by rendering regions empty of either any observations or observations for a given value of $L$. Three different pairs of $g_{A}(.)$ and $\sigma_{a}^{2}$ were examined:
@@
For $g_{Y}(.)$, which controls the shape of the causal dose-response curve (CDRC), a polynomial was used in order to provide a certain kind of complexity for the models to learn. As MTPs shift data points in the direction of treatment, the behaviour of this polynomial outside of the plotted interval of interest $[0,1]$ is relevant. In order to make the behaviour out of bounds simpler, the polynomial turns to a linear on the sides. The resulting data distribution and underlying truth, i. e., what the expected outcome would be if all observations received that treatment value, can be seen in Figure 3(a). The sample size $n$ was set to 350. Two other functional shapes for $Y$ where investigated as well, namely a piece-wise linear shape and a step function. The specifications for both of these are given in Appendix B. The full code for the simulation studies can be found in
Intervention Formula Dose Example (medication)
@@
naive shift-response curve (SRC) $\mathrm{E}\left(Y^{A=a+x}\right)$ $a_{i}+x\,\,\forall i$ increasing everyone’s dose by 5 g
@@
stochastic SRC $\mathrm{E}\left(Y^{A\sim\mathrm{N}(a+x,\sigma^{2})}\right)$ $a_{i}+x\pm\epsilon_{i}\,\,\forall i$ increasing everyone’s dose by 5 g on average
@@
threshold response curve $\mathrm{E}\left(Y^{A=\max(x,a)}\right)$$\left\{\begin{smallmatrix}x\hfill&\text{ if }a_{i}<x\hfill\\ &\\ a_{i}\hfill&\text{ else}\hfill\end{smallmatrix}\right.$ giving a dose of 5 g to subjects with a dose lower than 5 g (no change else)
@@
conditional shift response curve $\mathrm{E}\left(Y^{A=a+x\mathbbm{1}_{a+x\leq w}}\right)$$\left\{\begin{smallmatrix}a_{i}+x\hfill&\text{ if }a_{i}+x\leq w_{i}:=w(a_{i})\hfill\\ &\\ a_{i}\hfill&\text{ else}\hfill\end{smallmatrix}\right.$ increasing the dose by 5 g, but only if this increased dose is not toxic (no change else)
@@
Table 1: Listed are the different intervention curves considered for the single time point simulation study. The function $w$ is user defined based on the subject’s history $H$, which includes $A$. For the stochastic models, $\epsilon$ is drawn from $\mathrm{N}(\mu,\sigma^{2})$, where $\mu$ is provided and $\sigma^{2}$ estimated using a linear regression model.
@@
Table 1 shows the different intervention schemes we investigated in this simulation study. As static schemes are the most prevalent in research, we use the corresponding CDRC as a benchmark to illustrate how a switch to an MTP changes both estimand and estimability. The causal parameter of interest is here $\psi(P_{0})=\mathrm{E}\left(Y^{A=x}\right)$. The next scheme we considered is a naive shift and its naive shift-response curve. Here, we estimate $\psi(P_{0})=\mathrm{E}\left(Y^{A=a+x}\right)$. It is the simplest type of MTP. For the naive shift, we also considered its stochastic variant, namely $\psi(P_{0})=\mathrm{E}\left(Y^{A\sim\mathrm{N}(a+x,\sigma^{2})}\right)$ and $\mathrm{E}\left(Y^{A\sim\mathrm{N}(a+x,\sigma^{2})}\right)$, where the intervention scheme remains the same on average, but has a random component added. In addition to these rather simple MTPs we examined a threshold MTP and a shift MTP (in order to differentiate this scheme from the naive shift, we will refer to this as a conditional shift MTP) as the two highlighted schemes by Díaz et al. [8]. The threshold shift MTP only shifts observations below a certain threshold with estimate $\psi(P_{0})=\mathrm{E}\left(Y^{A=\max(x,a)}\right)$. The conditional shift MTP only shifts observations where a specific function is fulfilled, i. e., $\mathrm{E}\left(Y^{A=a+x\mathbbm{1}_{a+x\leq w}}\right)$. For the function $w(a)$, we simply used the maximum and minimum values (depending on the direction of the shift) occurring in the data for each given $L$, meaning no observations will be shifted to have a treatment value which exceeds the treatment range of the observed data of their $L$ group. We estimated shift-response curves for both of these. As we do not believe stochastic versions of either of these to be useful, we did not include them in the study.
@@
For each of the $3$ positivity situations, the dose-response curve was fitted using the CICI package with a g-formula based on a generalized GAM. Additionally, we used a doubly robust method based on Kennedy [2]. All MTP models were fitted with the R package lmtp’s Sequential Doubly Robust Estimator [11], as the authors recommended method, used two folds for cross-fitting, and the following algorithms in its Superlearners: glm, gam, glmnet, and earth. The hyperparameters of the diagnostic where set 0.15 as half distance for $A$. As three different data distributions are shown in this section, rather than using a different one based on the standard deviation for each scenario, a unified value was chosen to make comparisons easier.
@@ 4.2 Results
In this section we show the response curves for the different intervention schemes and compare estimates to the true response. The plots show how the true curves and the estimated curves differ and how the diagnostic could help provide the same information without needing access to the true response.
@@ 4.2.1 Causal Dose Response Curve
(a) dose response curves
@@
Figure 3(a) shows that in the overlapping data case there is a stretch of data with a lot of support, especially in the area where the density of $L=1$ is high. This is because $L=1$ is the rarer type of observation, meaning it represents a bottleneck. The functional shape of both the parametric g-formula estimate and the doubly robust estimate is very close to the truth. Both models are having trouble capturing the truth near the ends due to general sparsity. What is surprising is that around 0.25, where there is only data with $L=0$, both models still perform well. This is because the general trend is just continued from areas with more data. If this is not the case, for example if observations with $L=1$ had a value of 0.8 around this value, the estimates would be wrong. Basically, it was lucky that extrapolation worked well here, although this was not to be expected. The diagnostic reflects a conservative estimate on where estimation works well. Therefore only areas with large support for both types of $L$ are shown as well supported, all other areas have lower ends of the boxplots near zero. An interesting side note is, that the boxplots only represent two values as in a fixed scheme all $a_{i}$ are set to the same value so the only difference in observations is in $L$. The median is therefore the effective support of all observations with $L=0$ and the other end of the boxplot represents the effective points for all observations with $L=1$.
For the adjacent data structure the g-formula estimate performs well, but the doubly robust estimate is off, especially in areas with little data. The diagnostic shows that there is no region with good data support for all observations, as we intentionally induced a positivity issue. The parametric model performs well because its parametric assumptions are in line with the truth, the doubly robust estimator on the other hand does not benefit from any parametric assumption and does not perform well. Again, the diagnostic is a careful indicator and does not rely on any parametric assumptions being correct, which is why it deems the whole range as poorly supported.
Lastly, both estimators are far from the truth for the divided data. The diagnostic shows a support of approximately zero for the whole range, correctly reflecting a data situation in which estimation is not possible.
@@ 4.2.2 Shift Response Curve
In Figure 3(b) the naive shift response curve is plotted. Due to the differing data distributions in the three simulations, it first has to be noted that the true curves are different in each scenario. With a naive shift model, data are always shifted outside of the range and the correctness of the result depends on extrapolation working well into these areas without data. In the overlapping data case, it is clear this extrapolation works better in the negative shift direction. The diagnostic, however, shows that there are problems with support in both directions. In the adjacent data case, and even more in the divided data scenario, the data are more densely located, therefore smaller shifts mean the data range is left for more data points, which in turn means the estimation becomes more biased for smaller shifts. The more densely located data have another effect, too. For a shift of zero, the support is a lot better the more divided the data is.
@@ 4.2.3 Threshold Response Curve
For the threshold response curve in Figure 4(a) the diagnostic shows that support becomes a problem as soon as the red data points are shifted outside of the area with red data points. For very low thresholds (almost) no observations are intervened on, therefore the result is the natural intervention where every subject receives their observed treatment. This area gets smaller the more positivity is violated in this simulation. Therefore the model does not perform well for divided data, but is relatively stable for a large part of the threshold curve for the overlapping data. The diagnostic shows where the curve has support and can, without modeling, tell where the estimation will deviate from the true curve.
@@ 4.2.4 Conditional Shift Response Curve
For the conditional shift response curve in Figure 4(b) observations are just intervened on as long as they do not leave the range of their category, either $L=0$ or $L=1$. This means positivity issues are largely avoided. The diagnostic shows high support near zero and on the left and right side of the plots as there are no observations intervened on. The estimator does deviate from the the true curve in all scenarios, however it does not leave the confidence bounds. The diagnostic shows high support for the whole curve especially in the divided data case. Therefore, this estimand is a viable option in all three scenarios.
@@ 4.2.5 Further Simulations
The results for the simulations for the piece-wise linear shape and the step function are shown in Appendix B.
@@ 4.2.6 Conclusion
The results show that positivity violations cause problems for MTPs whenever observations get shifted into areas where extrapolation becomes unstable. This means it is vital to understand when a scheme shifts data into these areas. A diagnostic that shows us how much data support we have in order to estimate the effect of the intervention scheme for every observation in our dataset can help
@@ 5 Practical Application
In order to showcase how the diagnostic can be used by applied researchers, we apply it to data from a recent publication on determining targets for antiretroviral drug concentrations in children with HIV [18]. In this publication, the researchers evaluated the effect of a continuous exposure (drug concentration), measured over multiple time points, on a binary outcome (viral failure). We show how the diagnostic can be used to identify problems with positivity and other kinds of sparsity, and develop suggestions to combat these. For the example at hand, this could help develop an estimand that leads to more stable estimates than those considered in the study.
@@ 5.1 Data and Research Question
The data used for this study stems from the CHAPAS-3 (Children with HIV in Africa – Pharmacokinetics and Adherence/Acceptability of Simple antriretroviral regimens) trial, a randomized controlled trial, where HIV-positive children between 1 month and 13 years of age in Zambia and Uganda were given different combination antiretroviral therapy regimens [19]. The current study only uses the subset of participants who received the drug efavirenz, which leads to a sample size of 125 children. The authors conducted a complete case analysis, which they had previously shown to be valid for the question explored below given the missingness mechanism they assumed [20], leaving 58 children to be analyzed.
The target variable for this study was viral failure $VF$, defined as $>100$ copies/mL. The exposure variable was not the amount of medication taken, but the so-called mid-dose concentration of efavirenz $EFV$, meaning the concentration after 12 hours as medication is taken every 24 hours. This rather atypical exposure is utilized, because the target concentration window is at the core of the research question. This concentration was not directly measured, but calculated from measurements taken 2-4 hours after intake through a population PK model [21]. The data is longitudinal in nature and follow-up was at $t=6,\ 36,\ 48,\ 60,\ 84,\text{ and }96$ weeks. Viral load was measured at all time points except 6 and efavirenz concentrations were assessed all weeks except 48 and 96. Therefore the goal of the analysis is to determine $P(VF_{t}^{a})$ with $t=36,48,60,84,96$ and $a$ signifying a treatment scheme of efavirenz. This is in line with previous research on the same data set [22, 20, 18]. The sufficient adjustment set comprises only weight, here in a logarithmized form, and a time-independent indicator of general adherence to the treatment protocol [23]. Lastly, the metabolization group is of interest as a potential effect modifier as it is one of the main contributors to the high between-subject variability of efavirenz treatment [21]. The research question of the study concerns the CDRC for the antiretroviral drug efavirenz (more precisely in this case a Causal Concentration Response Curve) with a focus on the therapeutic window of the target concentration. In both adults and children the mid-dose target concentration window is typically given as between 1 mg/L and 4 mg/L, but especially in children the validity of this window is partially unclear [22, 18]. The upper end of the target window limits the probability of adverse events, but as these were rare in the trial, this upper limit cannot well be evaluated using the data at hand. Therefore, the focus lies on the lower bound, which limits the risk of viral failure.

### v3

@@ 4 Simulation Study
As previously stated, publications in the field of causal research are rather focused on static interventions. Other intervention schemes and MTPs are much less represented in research. Therefore, MTPs are not well established and generally not well understood by practitioners. However, the use of MTPs is often recommended as a response to positivity issues, as they “have the advantage that they can be designed to satisfy the positivity assumption required for causal inference” [9, p. 846]. In the following simulation study, we provide insights into the behaviour of some of the most popular MTPs in different sparsity situations. We show that it is not obvious when stochastic positivity issues cause bias given an estimand and demonstrate that our diagnostic bridges that gap, providing the information needed to decide which estimands are feasible and which are likely to be heavily biased due to sparsity issues. The code for this simulation study is made available in full on Github: https://github.com/KatyFisch/kernelbasedSparsityDiagnostic.
@@ 4.1 Methods
The simulation is kept as simple as possible in order to illustrate the issues MTPs face in light of positivity violations and how the diagnostic can help to understand the problem. It is meant as a proof of concept. An analysis of a more complex scenario is presented in the form of an applied study in Section 5. The data are simulated using the simple structural causal model $O=(L,A,Y)\sim P_{0}$, where $L$ is a binary confounder, $A$ the continuous treatment variable, and $Y$ the continuous outcome:
@@
The function $g_{A}(.)$ and variable $\sigma_{a}^{2}$ control the distribution of $A$ and can be used to induce problems regarding stochastic positivity by rendering regions empty of either any observations or specific observations for a given value of $L$. Three different pairs of $g_{A}(.)$ and $\sigma_{a}^{2}$ were examined:
@@
For $g_{Y}(.)$, which controls the shape of the causal dose-response curve (CDRC), a polynomial was used in order to provide a certain kind of complexity for the models to learn. As MTPs shift data points in the direction of treatment, the behaviour of this polynomial outside of the plotted interval of interest $[0,1]$ is relevant. In order to make the behaviour out of bounds simpler, the polynomial turns to a linear on the sides. The resulting data distribution and underlying truth, i. e., what the expected outcome would be if all observations received that treatment value, can be seen in Figure 3(a). The sample size $n$ was set to 350. Two other functional shapes for $Y$ were investigated as well, namely a piece-wise linear shape and a step function. The specifications for both of these are given in Appendix C.
Intervention Formula Dose $a_{i}^{int}$ Example (medication)
@@
naive shift-response curve (SRC) $\mathrm{E}\left(Y^{A=a^{obs}+x}\right)$ $a_{i}^{obs}+x\,\,\forall i$ increasing everyone’s dose by 5 g
@@
stochastic SRC $\mathrm{E}\left(Y^{A\sim\mathrm{N}(a^{obs}+x,\sigma^{2})}\right)$ $a_{i}^{obs}+x\pm\epsilon_{i}\,\,\forall i$ increasing everyone’s dose by 5 g on average
@@
threshold response curve $\mathrm{E}\left(Y^{A=\max(x,a^{obs})}\right)$ $\left\{\begin{smallmatrix}x\hfill&\text{ if }a_{i}^{obs}<x\hfill\\ &\\ a_{i}^{obs}\hfill&\text{ else}\hfill\end{smallmatrix}\right$. giving a dose of 5 g to subjects with a dose lower than 5 g (no change else)
@@
conditional shift response curve $\mathrm{E}\left(Y^{A=a^{obs}+x\mathbbm{1}_{a^{obs}+x\leq w}}\right)$ $\left\{\begin{smallmatrix}a_{i}^{obs}+x\hfill&\text{ if }a_{i}^{obs}+x\leq w_{i}\hfill\\ &\\ a_{i}^{obs}\hfill&\text{ else}\hfill\end{smallmatrix}\right$. increasing the dose by 5 g, but only if this increased dose is not toxic (else no change)
@@
Table 1: List of the different intervention curves considered for the single time point simulation study. The function $w$ is user defined based on the subject’s history $H$, which includes AA. For the stochastic models, $\epsilon$ is drawn from $\mathrm{N}(\mu,\sigma^{2})$, where $\mu$ is provided and $\sigma^{2}$ estimated using a linear regression model.
@@
Table 1 shows the different intervention schemes we investigated in this simulation study. As static schemes are the most prevalent in research, we use the corresponding CDRC as a benchmark to illustrate how a switch to an MTP changes both estimand and estimability. The causal parameter of interest is here $\psi(P_{0})=\mathrm{E}\left(Y^{A=x}\right)$. The next scheme we considered is a naive shift and its naive shift-response curve. Here, we estimate $\psi(P_{0})=\mathrm{E}\left(Y^{A=a^{obs}+x}\right)$. It is the simplest type of MTP. For the naive shift, we also considered its stochastic variant, namely $\psi(P_{0})=\mathrm{E}\left(Y^{A\sim\mathrm{N}(a^{obs}+x,\sigma^{2})}\right)$, where the intervention scheme remains the same on average, but has a random component added. In addition to these rather simple MTPs, we examined a threshold MTP and a shift MTP (in order to differentiate this scheme from the naive shift, we will refer to this as a conditional shift MTP), which coincide with the two schemes highlighted by Díaz et al. [9]. The threshold shift MTP only shifts observations below a certain threshold with estimate $\psi(P_{0})=\mathrm{E}\left(Y^{A=\max(x,a^{obs})}\right)$. The conditional shift MTP only shifts observations where a specific function is below a certain threshold, i. e., $\mathrm{E}\left(Y^{A=a^{obs}+x\mathbbm{1}_{a+x\leq w}}\right)$. For the function $w_{i}=w(a_{i}^{obs})$, we simply used the maximum and minimum values (depending on the direction of the shift) occurring in the data for each given $L$, meaning no observations will be shifted to have a treatment value which exceeds the treatment range of the observed data of their $L$ group. We estimated shift-response curves for both of these. As we do not believe stochastic versions of either of these to be useful, we did not include them in the study.
@@
For each of the $3$ stochastic positivity situations, the dose-response curve is fitted using the CICI package with a g-formula based on a generalized GAM. Additionally, we use a doubly robust method based on Kennedy [2]. All MTP models are fitted with the R package lmtp’s Sequential Doubly Robust Estimator [15], which is the authors’ recommended method, use two folds for cross-fitting, and the following algorithms in its Superlearners: glm, gam, glmnet, and earth. The hyperparameters of the diagnostic were set to 0.15 as half distance for $A$. As three different data distributions are shown in this section, rather than using a different one based on the standard deviation for each scenario, a unified value was chosen to make comparisons easier.
@@ 4.2 Results
In this section we show the response curves for the different intervention schemes and compare estimates to the true response. The plots show how the true curves and the estimated curves differ and how the diagnostic could help provide the information where bias is to be expected without needing access to the true response.
@@ 4.2.1 Causal Dose Response Curve
(a) dose response curve
@@
Figure 3(a) shows that in the overlapping data case there is a stretch of data with a lot of support for both $L=1$ and $L=0$, especially in the area where the density of $L=1$ is high. This is because $L=1$ is the rarer type of observation, meaning it represents a bottleneck. The functional shape of both the parametric g-formula estimate and the doubly robust estimate is very close to the truth. Both models are having trouble capturing the truth near the ends due to general sparsity. What is surprising is that around 0.25, where there is only data with $L=0$, both models still perform well. This is because the general trend is just continued from areas with more data. If this is not the case, for example if observations with $L=1$ had a value of 0.8 around this value, the estimates would be wrong. Basically, it is due to luck that extrapolation worked well here, although this was not to be expected. The diagnostic reflects a conservative estimate on where estimation works well. Therefore only areas with large support for both types of $L$ are shown as well supported, all other areas have lower ends of the boxplots near zero. An interesting side note is, that the boxplots only represent two values as in a fixed scheme all $a_{i}$ are set to the same value so the only difference in observations is in $L$. The median is therefore the effective support of all observations with $L=0$ and the other end of the boxplot represents the effective points for all observations with $L=1$.
For the adjacent data structure the g-formula estimate performs well, but the doubly robust estimate is off, especially in areas with little data. The diagnostic shows that there is no region with good data support for all observations, as we intentionally induced a stochastic positivity issue. The parametric model performs well because its parametric assumptions are in line with the truth, the doubly robust estimator on the other hand does not benefit from any parametric assumption and does not perform well. Again, the diagnostic is a careful indicator and does not rely on any parametric assumptions being correct, which is why it deems the whole range as poorly supported.
Lastly, both estimators are far from the truth for the divided data. The diagnostic shows a support of approximately zero for many data points for the whole range, correctly reflecting a data situation in which satisfactory estimation is not possible.
@@ 4.2.2 Shift Response Curve
In Figure 3(b) the naive shift response curve is plotted. Due to the differing data distributions in the three simulations, it first has to be noted that the true curves are different in each scenario. With a naive shift model, some data points are always shifted outside of the range and the correctness of the result depends on extrapolation working well into these areas without data. In the overlapping data case, it is clear this extrapolation works better in the negative shift direction. The diagnostic, however, shows that there are problems with support in both directions. In the adjacent data case, and even more in the divided data scenario, the data points are more densely situated, therefore smaller shifts mean more data points leave the data range, which in turn means the estimation becomes more biased for smaller shifts. The more densely located data have another effect, too. In case of a shift of zero, the support is a lot better the more divided the data is.
@@ 4.2.3 Threshold Response Curve
For the threshold response curve in Figure 4(a) the diagnostic shows that support becomes a problem as soon as the green data points are shifted outside of the area spanned by the green observed data points. For very low thresholds (almost) no observations are intervened on, therefore the result is the natural intervention where every subject receives their observed treatment. This area is smaller the more stochastic positivity is violated in this simulation. Therefore the model does not perform well for divided data, but is relatively stable for a large part of the threshold curve in case of the overlapping data. The diagnostic shows where the curve has support and can, without modeling, tell where the estimation will deviate from the true curve.
@@ 4.2.4 Conditional Shift Response Curve
Concerning the conditional shift response curve in Figure 4(b), observations are intervened on as long as they do not leave the observed range of their category, either $L=0$ or $L=1$. This means, positivity issues are largely avoided. The diagnostic shows high support near zero and on the left and right side of the plots as there are no observations intervened on. The estimator does deviate from the true curve in all scenarios, however it does not leave the confidence bounds. The diagnostic shows high support for the whole curve especially in the divided data case. Therefore, this estimand is a viable option in all three scenarios.
@@ 4.2.5 Further Simulations
The results for the simulations of the piece-wise linear shape and the step function are shown in Appendix C. These additional simulations especially highlight how outliers, for which a prediction cannot be estimated well, could potentially lead to a strong bias in effect estimation, if the true outcomes for these outliers are extreme. They also shine a light on true response curves of non-polynomial form and therefore demonstrate that the diagnostic works in various data scenarios.
@@ 4.2.6 Conclusion
The results show that stochastic positivity violations cause problems for MTPs whenever observations get shifted into areas where extrapolation becomes unstable. This means it is vital to understand when a scheme shifts data into these areas. A diagnostic that shows how much data support there is in order to estimate the effect of the intervention scheme for every observation in the dataset can help get deeper insights.
@@ 5 Practical Application
In order to showcase how the diagnostic can be used by applied researchers, we apply it to data from a recent publication on determining targets for antiretroviral drug concentrations in children with HIV [21]. In this publication, the researchers evaluated the effect of a continuous exposure (drug concentration), measured over multiple time points, on a binary outcome (viral failure). We show how the diagnostic can be used to identify problems with stochastic positivity and other kinds of sparsity, and develop suggestions to combat these. For the example at hand, this could help develop an estimand that leads to more stable estimates than those considered in the study.
@@ 5.1 Data and Research Question
The data used for this study stems from the CHAPAS-3 (Children with HIV in Africa – Pharmacokinetics and Adherence/Acceptability of Simple antiretroviral regimens) trial, a randomized controlled trial, where HIV-positive children between 1 month and 13 years of age in Zambia and Uganda were given different combination antiretroviral therapy regimens [22]. The current study only uses the subset of participants who received the drug efavirenz, which leads to a sample size of 125 children. The authors conducted a complete case analysis, which they had previously shown to be valid for the question explored below given the missingness mechanism they assumed [23], leaving 58 children to be analyzed.
[…]